OpenStudy (sanra123):

algebra question please help: sam buys 10 t shirts and lilly buys 4 sweatshirts both pay the same amount but both have different prices and the total of both 10 t shirts and 4 sweatshirts is 230$ what is the price of each sweatshirt?

OpenStudy (thatonegirl_):

okay so first you are going to want to make equations with the info you know. You will have 2 equations. Any ideas on 1?

OpenStudy (sanra123):

is it 230+10x?

OpenStudy (sanra123):

ugh no thats wrong no i really don't know

OpenStudy (thatonegirl_):

Okay well we can figure out a price equation and a number equation (number of shirts)

OpenStudy (sanra123):

is it right?

OpenStudy (thatonegirl_):

so first we can pick 2 variables for sweatshirts and tshirts. how does s and t sound?

OpenStudy (thatonegirl_):

Not exactly.

OpenStudy (sanra123):

great)

OpenStudy (thatonegirl_):

Okay cool. We can start with the cost equation first since that one tends to be easier. We know from the info given that 10 tshirts and 4 sweatshirts cost $230 right?

OpenStudy (sanra123):

yes

OpenStudy (thatonegirl_):

So we can make an equation using those numbers and using the variables t and s which will represent the cost of 1 tshirt or 1 sweatshirt (depending on which variable it is). 10t+4s=230.

OpenStudy (sanra123):

okay im with u

OpenStudy (sanra123):

u know i tried this but then i didnt know what to do

OpenStudy (thatonegirl_):

okay good :) does that all make sense why you would set it up that way?

OpenStudy (sanra123):

yes

OpenStudy (thatonegirl_):

Lol well its good you tried! We all get stuck sometimes :/

OpenStudy (sanra123):

yes i know :(

OpenStudy (thatonegirl_):

Okay so the other info we have left is that 10 tshirts and 4 sweatshirts cost the same, right?

OpenStudy (sanra123):

not each one costs the same but when they purchased 10 t shirts it was the same as 4 sweatshirts

OpenStudy (thatonegirl_):

the total of 10 tshirts is the same total of 4 sweatshirts (meaning sweatshirts will be more expensive if you think about it. You get less sweatshirts than tshirts but it all totals up to the same amount)

OpenStudy (sanra123):

ye s yes

OpenStudy (sanra123):

yes yes thats true

OpenStudy (thatonegirl_):

Yup! That's a good catch. So when we write the equation we will need to make sure that we include 10 and 4 and not just 1 like you said.

OpenStudy (sanra123):

yes

OpenStudy (thatonegirl_):

So if they both cost the same amount, you could set them equal right? Does this part make sense?

OpenStudy (thatonegirl_):

10t=4s

OpenStudy (sanra123):

yes

OpenStudy (thatonegirl_):

Because the 10t is the COST, not the number of shirts. and 4s is the COST also (since t and s are prices per one shirt). And we know the costs are EQUAL. So therefore we can write that equation, good?

OpenStudy (sanra123):

yes yes

OpenStudy (thatonegirl_):

awesome. Here are our 2 equations: \[10t+4s=230\] \[10t=4s\]

OpenStudy (sanra123):

yes great :)

OpenStudy (thatonegirl_):

Any ideas on what we do with this excellent info we have? :D

OpenStudy (sanra123):

no idea :(

OpenStudy (sanra123):

but i understand everything till now

OpenStudy (thatonegirl_):

Okay don't worry :) The next goal is to solve for a variable! You can pick which you want to solve for first, t or s. We will eventually find both anyways. Which would you like to solve for first?

OpenStudy (sanra123):

t

OpenStudy (thatonegirl_):

Okay. So then we will need to get only t in an equation. 10t+4s=230 <--- we can use this equation to solve for t. (you could use the other but this is one is easier in my opinion) 10t=4s <--- here we have it telling us that 4s=10t. That means we can do... \[10s+10s=230\] (this is the first equation just with the substitution)

OpenStudy (thatonegirl_):

Do you see how i got that?

OpenStudy (sanra123):

whattt

OpenStudy (sanra123):

why 10s +10s

OpenStudy (thatonegirl_):

oooh my bad, I meant 10t+10t!!

OpenStudy (thatonegirl_):

not s lol

OpenStudy (thatonegirl_):

Does that clear it up, or do you want me to explain still?

OpenStudy (sanra123):

wait why did we do this equation ugh sorry

OpenStudy (sanra123):

how did u get it

OpenStudy (thatonegirl_):

10t+4s=230 10t=4s This is the info we have. 10t+*4s*=230 we need to get only t's in the equation, BUT there is a sneaky *4s* in there. It's no biggie though because in our other equation, it tells us that 10t=*4s* which is the same as *4s*=10t. If 4s is the same thing as 10t then you can plug in 10t for 4s! Because they equal the same thing, so the value is the same, just a different way of writing it :D It won't change the equations value, just different variables, but it's still equal :)

OpenStudy (thatonegirl_):

10t+10t=230

OpenStudy (sanra123):

ohhhhh okay so basically u subistusted because they are equal

OpenStudy (thatonegirl_):

Exactly!!

OpenStudy (thatonegirl_):

Now you can solve for t because it's the only variable in the equation. Let me know what you get

OpenStudy (sanra123):

so its 20t=230

OpenStudy (sanra123):

so its 11.5

OpenStudy (thatonegirl_):

Perfect, good job! :D So if t (price of 1 tshirt) is $11.50, we can plug that info into either equation for t and find the price of a sweatshirt (s).

OpenStudy (thatonegirl_):

I should've reread your question, it only wants to know the price of a sweatshirt, not the tshirt. But more info is never a bad thing :P

OpenStudy (sanra123):

yes hahaha i was going to tell u buts its my bad

OpenStudy (thatonegirl_):

10t+4s=230 10(11.5)+4s=230 OR you can do this one 10t=4s 10(11.5)=4s

OpenStudy (thatonegirl_):

Lol no problem xD Extra practice is good

OpenStudy (sanra123):

let me solve it and tell u the answer

OpenStudy (thatonegirl_):

You can pick whichever one you want, you don't HAVE to solve for both, they should give the same answer

OpenStudy (sanra123):

28.75

OpenStudy (sanra123):

waiiitt u know i did it another way without doing equations and i had the same answer

OpenStudy (thatonegirl_):

You got it!! And how did you do it?

OpenStudy (sanra123):

i divided 230/2 since they are equal then i got 115 then i did 115/10 which is 11.5 and 115/4 which is 28.7

OpenStudy (thatonegirl_):

Oh wow, that is a much easier and faster way and it definitely works! Super smart of you!!

OpenStudy (sanra123):

yes but i think your way is better since its algebra and we need formulas there u know

OpenStudy (sanra123):

your so smart i really like how u explain everything really detailed so smart way to go !

OpenStudy (thatonegirl_):

Yes sadly :C But knowing multiple ways of doing it is always a good thing, so great job!

OpenStudy (thatonegirl_):

Aww thank you! I hope it helped :)

OpenStudy (sanra123):

thank you and i really appreciate your help i really wanted someone to help and u were there so i really appreciate that

OpenStudy (thatonegirl_):

Anytime!!! Good luck with the rest of your work :D

OpenStudy (sanra123):

it helped a lot thank you again :)

OpenStudy (sanra123):

thank you